### v1

@@ 3 Correlated Gaussian Mechanism for DP-DME: A Geometric Interpretation
In this section, we provide a geometric interpretation of the DP-DME problem, where we show how carefully tuned (anti) correlated noise can significantly outperform DME with independent noise (LDP) while ensuring the same level of privacy, in the absence of dropouts and collusions. This also presents a re-interpretation of one of the core ideas of the SecAgg scheme - but over $\mathbb{R}^{d}$ in the DP framework rather than over finite fields. Building up on this idea, we analyze the privacy-utility trade-offs with the presence of dropouts and/or collusions, and show that carefully tuned correlated noise continues to provide superior performance to independent noise (see Section 5 for the results).
@@
Consider a simple two-user setting with no dropouts and no colluding users, where the private vectors of the two users are given by $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, with $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{S}^{d-1}$. User $i$ sends $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$ to the server, where $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d})$ for $i=1,2$. The elements of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ are correlated as,
$\mathbb{E}[Z_{1,j}Z_{2,k}]=\begin{cases}\rho\sigma^{2},&j=k,\quad j,k\in\{1,\dotsc,d\}\\ 0,&j\neq k,\quad j,k\in\{1,\dotsc,d\}.\end{cases}$ (7)
@@
Next, we geometrically illustrate the privacy constraint in (3). For this example, (3) simplifies to $\mathbb{P}(\mathbf{x}_{i}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{Z}_{j}=\mathbf{y}_{j}-\mathbf{x}_{j})\leq e^{\epsilon}\mathbb{P}(\mathbf{x}_{i}^{\prime}+\mathbf{Z}_{i}\in\mathcal{A}|\mathbf{Z}_{j}=\mathbf{y}_{j}-\mathbf{x}_{j})+\delta$, for each $i\neq j$, for any fixed $\mathbf{x}_{j},\mathbf{y}_{j}$ and $\forall\mathbf{x}_{i},\mathbf{x}_{i}^{\prime}$. Let $\bm{Z}_{i}^{\parallel}:=\langle\bm{Z}_{i},\bm{Z}_{j}\rangle\bm{Z}_{j}/|\bm{Z}_{j}|^{2}=(\cos\theta)\bm{Z}_{j}$ and $\bm{Z}_{i}^{\bot}:=\bm{Z}_{i}-\bm{Z}_{j}^{\parallel}$ be the components of $\bm{Z}_{i}$ in $\mathcal{H}$, that are parallel and orthogonal to $\bm{Z}_{j}$, respectively, for $j\neq i$. The privacy constraint on $\bm{x}_{i}$ essentially imposes the lower bound: $|\bm{Z}_{i}^{\bot}|\geq\hat{\sigma}_{\epsilon,\delta}$, where $\hat{\sigma}_{\epsilon,\delta}$ is a constant that depends on the given privacy parameters and ||$\cdot$|| ⋅ | denotes the magnitude of a vector in $\mathcal{H}$ (See Appendix F for a rigorous proof). This is illustrated in Fig. 2, along with the vectors $\bm{Z}_{1}+\bm{Z}_{2}$ in red, that represent the MSEs corresponding to 1) $\rho<0$, 2) $\rho=0$, and 3) $\rho>0$. As we assume Gaussian noise, $\rho=0$ corresponds to independent noise, and hence translates to classical LDP. If $\rho=0,$ then the minimum MSE is $|\bm{Z}_{1}+\bm{Z}_{2}|^{2}=E[||\bm{Z}_{1}+\bm{Z}_{2}||^{2}]=2\hat{\sigma}_{\epsilon,\delta}^{2}$ (this can also be observed geometrically from Fig. 2). However, notice that when $\rho<0$, certain components of $\bm{Z}_{1}$ and $\bm{Z}_{2}$ cancel out and results in a reduced $|\bm{Z}_{1}+\bm{Z}_{2}|$, compared to the other two cases. In the context of Fig. 2, it is required to optimize $|\bm{Z}_{i}|$ and $\theta$, i.e., $\sigma^{2}$ and $\rho$, to minimize $|\bm{Z}_{1}+\bm{Z}_{2}|$ such that $|\bm{Z}_{i}^{\bot}|=|\bm{Z}_{i}|\sin{\theta}\geq\hat{\sigma}_{\epsilon,\delta}$ for $i=1,2$. The optimum $\sigma^{2}$ and $\rho$ are given by $\sigma^{2}_{*}\to\infty$ and $\rho_{*}\to-1$. This result is explained as follows. As illustrated in Fig. 3, $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|$ decreases as $|\mathbf{Z}_{i}|$ and $\theta$ increase while satisfying $|\mathbf{Z}_{i}|\sin\theta=\hat{\sigma}_{\epsilon,\delta}$ for privacy. In the limit, when $|\mathbf{Z}_{i}|\to\infty$ and $\theta\to\pi$,***Note that increasing $\theta$ beyond $\pi$ is not optimal as $\theta=\pi+\eta$ and $\theta=\pi-\eta$ correspond to the same setting. $\mathbf{Z}_{1}+\mathbf{Z}_{2}$ aligns perpendicular to $\mathbf{Z}_{2}$, and $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|\to\hat{\sigma}_{\epsilon,\delta}$, which is the minimum achievable $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|$ while ensuring $|\mathbf{Z}_{i}|\sin\theta=\hat{\sigma}_{\epsilon,\delta}$.The resulting minimum MSE is given by $\mathbb{E}[\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|^{2}]=|\mathbf{Z}_{1}+\mathbf{Z}_{2}|^{2}=\hat{\sigma}^{2}_{\epsilon,\delta}$.
Figure 2: The privacy constraint on $\mathbf{x}_{1}$, described by a lower bound on $|\mathbf{Z}_{1}^{\bot}|$, and the MSE, represented by $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|$.
Figure 3: Variation of $|\mathbf{Z}_{1}+\mathbf{Z}_{2}|$ with changing $|\mathbf{Z}_{i}|$.
Next, we analyze the value of $\hat{\sigma}_{\epsilon,\delta}$ for any given $\epsilon$ and $\delta$. Let $\sigma_{\epsilon,\delta}=\inf_{\hat{\sigma}>0}\{\hat{\sigma};\Phi\left(\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)-e^{\epsilon}\Phi\left(-\frac{1}{\hat{\sigma}}-\frac{\epsilon\hat{\sigma}}{2}\right)\leq\delta\}$ for any given $\epsilon$ and $\delta$. The standard Gaussian mechanism in DP states that $\mathbf{Z}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d})$ with $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ ensures $\mathbb{P}(\mathbf{v}+\mathbf{Z}\in\mathcal{A})\leq e^{\epsilon}\mathbb{P}(\mathbf{v}^{\prime}+\mathbf{Z}\in\mathcal{A})+\delta$ for any $\mathbf{v},\mathbf{v}^{\prime}\in\mathbb{R}^{d}$ with $\|\mathbf{v}-\mathbf{v}^{\prime}\|\leq 2$, and $\forall\mathcal{A}\subset\mathbb{R}^{d}$. Note that this is the setting of CDP where $\mathbf{v}$ and $\mathbf{v}^{\prime}$ represent the summations of the users’ vectors with all vectors fixed except for one, which results in $\|\mathbf{v}-\mathbf{v}^{\prime}\|\leq 2$ as each vector is placed within the unit ball. Equivalently, for the case of CDP, the estimate of $\mathbf{v}$ given by $\mathbf{v}+\mathbf{Z}$ incurs an MSE that is lower bounded by,
$\mathbb{E}[\|\mathbf{Z}\|^{2}]\geq\sigma^{2}_{\epsilon,\delta}d$ (8)
to satisfy $(\epsilon,\delta)$-DP. Recall that when considering the privacy of $\mathbf{x}_{i}$ in the two user case, the effective noise added to $\mathbf{x}_{i}$ is quantified by $\mathbf{Z}_{i}^{\bot}$ as the rest of $\mathbf{Z}_{i}$ can be inferred by $\mathbf{Z}_{j}$, for $i,j\in\{1,2\}$, $j\neq i$. The privacy constraint in (3) reduces to $\mathbb{P}(\mathbf{x}_{i}+\mathbf{Z}_{i}^{\bot}\in\mathcal{A}^{\prime})\leq e^{\epsilon}\mathbb{P}(\mathbf{x}^{\prime}_{i}+\mathbf{Z}_{i}^{\bot}\in\mathcal{A}^{\prime})+\delta,\quad i=\{1,2\}$ for $i=\{1,2\}$, $\forall\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{B}^{d}$, and $\forall\mathcal{A}^{\prime}\subset\mathbb{R}^{d}$ in the Borel $\sigma$-algebra. This imposes a lower bound on the variance of $\mathbf{Z}_{i}^{\bot}$, and hence on the MSE resulted by the effective noise: $\mathbb{E}[\|\mathbf{Z}_{i}^{\bot}\|^{2}]=|\mathbf{Z}_{i}^{\bot}|^{2}\geq\sigma^{2}_{\epsilon,\delta}d$ to ensure $(\epsilon,\delta)$-DP based on the standard Gaussian mechanism, similar to (8). This implies $\hat{\sigma}_{\epsilon,\delta}=\sigma_{\epsilon,\delta}\sqrt{d}$.
Recall from (8) that the minimum MSE of DME with CDP is given by $\sigma^{2}_{\epsilon,\delta}d=\hat{\sigma}^{2}_{\epsilon,\delta}$. This (two-user) example shows that the same MSE can be achieved without the requirement of a trusted server by carefully choosing the parameters of the correlated privacy mechanism; notably, this is a factor of $2$ reduction as compared to LDP ($\rho=0$). The insights of Fig. 2 generalize for more than two users. Specifically, in Theorem 1 and Corollary 1, we show that the same MSE of $\sigma^{2}_{\epsilon,\delta}d$ is achieved for the general case of $n\geq 2$ users with carefully correlated privacy mechanisms, even with no trusted server. If the privacy mechanisms are not correlated, i.e., if $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d})$ are independent for $i\in[1:n]$, $\sigma^{2}\geq\sigma^{2}_{\epsilon,\delta}$ must hold for privacy and the estimation error is $\sum_{i=1}^{n}\mathbf{Z}_{i}$, which results in a minimum MSE of $\mathbb{E}[\|\sum_{i=1}^{n}\mathbf{Z}_{i}\|^{2}]=n\sigma^{2}_{\epsilon,\delta}d$, which is $n$ times larger than the MSE achieved by DME with CorDP-DME for the same level of privacy.
@@
The correlated Gaussian scheme described above faces a significant shortcoming. If a user drops out, the residual noise of the remaining user/s cannot be canceled resulting in significantly high MSEs; this is because it uses noise with arbitrarily large variance (recall that the lowest possible MSE is achieved when $|\mathbf{Z}_{i}|^{2}=E[||\mathbf{Z}_{i}||^{2}]\to\infty$). In fact, SecAgg faces the same shortcoming. In SecAgg, users’ data is quantized to a finite field, and then a noise random variable that is uniformly distributed over the field elements is added to the quantized data and sent to the server. If a user drops out, this added noise cannot be canceled and the server input from the remaining users is statistically independent of the users’ data - which in effect results in a large MSE. SecAgg circumvents dropouts through additional rounds where remaining users’ share these non-canceled noise variables with the server enabling the server to cancel it***In fact, SecAgg is complex and requires even an additional third round in order it has to guard against data leaks and malicious behaviour in this second round as well.. However, our interpretation here opens the door to handling user dropouts much more directly, leading CorDP-DME protocol. In CorDP-DME we optimize the noise parameters—$\rho$ and $\sigma^{2}$—directly accounting for dropouts (up to a threshold). As our analysis demonstrates, this approach significantly improves upon LDP in terms of MSE, while avoiding additional rounds of communication. In fact, in Section 5, we show that for any dropout threshold, CorDP-DME outperforms LDP with independent Gaussian noise.
4 Proposed Approach

### v2

@@ III Correlated Gaussian Mechanism for DP-DME: A Geometric Interpretation
In this section, we illustrate the privacy-utility trade-off of DP-DME using a geometric approach. We show how carefully tuned (anti) correlated noise can significantly outperform DME with independent noise (LDP) while ensuring the same level of privacy, with no dropouts. This also presents a re-interpretation of the SecAgg scheme — but over $\mathbb{R}^{d}$ in the DP framework rather than over finite fields. Building up on this idea, in the subsequent sections of this paper, we analyze the privacy-utility trade-offs even with the presence of dropouts and/or collusions, and show that carefully tuned correlated noise continues to provide superior performance to independent noise (see Section IV-A for the results).
Figure 3: Geometric interpretation of the privacy-utility trade-off in DP-DME for different correlation coefficients among the users’ noise distributions: Noise vectors $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ are represented as vectors in $\mathcal{H}$ with magnitude $\sigma\sqrt{d}$ and angle $\theta=\cos^{-1}\rho$ between them. The privacy constraint on $\mathbf{x}_{i}$ enforces that the orthogonal component of $\mathbf{Z}_{i}$ relative to $\mathbf{Z}j$ (for $i\neq j$) has a magnitude bounded below by a constant $\gamma_{\epsilon,\delta}$. The MSE is proportional to the magnitude of $\mathbf{Z}_{1}+\mathbf{Z}_{2}$.
@@
Consider a simple two-user setting with no dropouts and no colluding users, where the private vectors of the two users are given by $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, with $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathbb{B}^{d}$. User $i$ sends $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$ to the server, where $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d})$ for $i=1,2$, with $\mathbf{0}_{d}$, $\mathsf{I}_{d}$, and $\sigma^{2}$ representing the all zeros vector of size $d\times 1$, identity matrix of size $d\times d$, and a constant, respectively. The elements of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ are correlated as follows:
$\mathbb{E}[Z_{1,j}Z_{2,k}]=\begin{cases}\rho\sigma^{2},&j=k,\quad j,k\in\{1,\dotsc,d\}\\ 0,&j\neq k,\quad j,k\in\{1,\dotsc,d\},\end{cases}$ (8)
where $Z_{i,j}$ is the $j$-th element of $\mathbf{Z}_{i}$ for $i\in\{1,2\}$, $j\in\{1,\dotsc,d\}$ and $\rho$ is the correlation coefficient between $Z_{1,k}$ and $Z_{2,k}$. For simplicity, let the decoder be $\boldsymbol{\alpha}$=[1,1]. Therefore, the server’s estimation of the mean of the vectors of the two users is given by $\hat{S}=\frac{1}{2}(\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{Z}_{1}+\mathbf{Z}_{2})$. The estimation error is $\frac{1}{2}(\mathbf{Z}_{1}+\mathbf{Z}_{2})$, which is quantified in terms of the MSE given by $\frac{1}{4}\mathbb{E}[\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|^{2}]$. The goal is to minimize $\mathbb{E}[\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|^{2}]$ while satisfying the privacy constraint in (3).
To define the problem geometrically, consider the following vector representation. Let $\mathcal{H}$ be an inner product space consisting of all Gaussian random variables of dimension $d$. For any $A,B\in\mathcal{H}$, let the inner product be defined as $\langle A,B\rangle_{\mathcal{H}}=\mathbb{E}[A^{T}B]$. With this definition, the noise vectors $\mathbf{Z}_{1},\mathbf{Z}_{2}$ in this example are represented as vectors with magnitude:
$\|\mathbf{Z}_{i}\|_{\mathcal{H}}=\sqrt{\langle\mathbf{Z}_{i},\mathbf{Z}_{i}\rangle_{\mathcal{H}}}=\sqrt{\mathbb{E}[\mathbf{Z}_{i}^{T}\mathbf{Z}_{i}]}=\sqrt{d\sigma^{2}}.$ (9)
The angle between $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ in $\mathcal{H}$ (denoted by $\theta$) corresponds to the correlation coefficient $\rho$ as:
$\langle\mathbf{Z}_{1},\mathbf{Z}_{2}\rangle_{\mathcal{H}}=d\sigma^{2}\cos\theta=\mathbb{E}[\mathbf{Z}_{1}^{T}\mathbf{Z}_{2}]=\rho\sigma^{2}d$ (10)
@@
Next, we illustrate the privacy constraint. For this example, (3) simplifies to,
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A}|M(\mathbf{x}_{j}))\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|M(\mathbf{x}_{j}))+\delta,$ (11)
for $i,j\in\{1,2\}$, $i\neq j$. To geometrically illustrate the privacy constraint, define,
$\boldsymbol{Z}_{i}^{\parallel}:=\frac{\langle\boldsymbol{Z}_{i},\boldsymbol{Z}_{j}\rangle_{\mathcal{H}}\boldsymbol{Z}_{j}}{\|\boldsymbol{Z}_{j}\|_{\mathcal{H}}^{2}}=(\cos\theta)\boldsymbol{Z}_{j}$ (12)
and $\boldsymbol{Z}_{i}^{\bot}:=\boldsymbol{Z}_{i}-\boldsymbol{Z}_{j}^{\parallel}$ to be the components of $\boldsymbol{Z}_{i}$ in $\mathcal{H}$, that are parallel and orthogonal to $\boldsymbol{Z}_{j}$, respectively, for $j\neq i$. The privacy constraint on $\boldsymbol{x}_{i}$ in (11) simplifies to:
$\|\boldsymbol{Z}_{i}^{\bot}\|_{\mathcal{H}}\geq\gamma_{\epsilon,\delta},\quad i\in\{1,2\}$ (13)
where $\gamma_{\epsilon,\delta}$ is a constant that depends on the given $\epsilon$ and $\delta$. The intuition behind this constraint is as follows. $\mathbf{Z}_{i}^{\bot}$ denotes the component of $\mathbf{Z}_{i}$ that remains independent of $\mathbf{Z}_{j}$. The uncertainty in $\mathbf{Z}_{i}^{\bot}$ is critical for preserving the privacy of $\mathbf{x}_{i}$, thus requiring a lower bound on its variance (See Appendix F-A for a rigorous proof).
Our goal of minimizing MSE while satisfying the privacy constraint is equivalent to minimizing $\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|_{\mathcal{H}}^{2}$ (red vectors in Fig. 3) while preserving $\|\boldsymbol{Z}_{i}^{\bot}\|_{\mathcal{H}}\geq\gamma_{\epsilon,\delta}$ (green dotted line in Fig. 3). The core insight of CorDP-DME is to find the best $\sigma$ (the norm of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$) and $\rho$ (the angle between $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$).
To gain insight into how $\rho$ and $\sigma$ affect the MSE for a given level of privacy, let us start with the simplest case where $\rho=0$. Since we assume Gaussian noise, $\rho=0$ corresponds to independent noise. Hence, this is simply LDP: adding sufficient noise at each user by setting $\|\boldsymbol{Z}_{i}\|_{\mathcal{H}}\geq\gamma_{\epsilon,\delta}$, without any correlation in the noise. Due to this lack of correlation, the angle between the noise vectors is always a right angle. As a result, $\|\boldsymbol{Z}_{1}+\boldsymbol{Z}_{2}\|_{\mathcal{H}}^{2}=2\cdot d\sigma^{2}$. When there are more than two users ($n>2$), it is straightforward to see that the total noise will be $n\cdot d\sigma^{2}$.
Now, let us consider the case when $\rho\neq 0$. When $\rho>0$ (the leftmost plot in Fig.3), this results in $\theta<\pi/2$. At the same time, to maintain $\|\boldsymbol{Z}_{i}\|_{\mathcal{H}}\geq\gamma_{\epsilon,\delta}$, the norms of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ must increase. These two factors together increase $\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|_{\mathcal{H}}^{2}$, which is undesirable. However, when $\rho<0$ (the rightmost plot in Fig.3), which corresponds to $\theta>\pi/2$, the anti-correlated components of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ cancel out, making $\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|_{\mathcal{H}}^{2}$ smaller than when $\rho=0$. Note that $\sigma$ still has to grow in this case to meet the privacy requirement.
As we understand how anti-correlated noise can reduce MSE while satisfying the same privacy constraint, let us consider how best we can design the anti-correlation. As you can see in Fig. 4, as $\theta$ increases, $\sigma$ must also grow, but $\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|_{\mathcal{H}}^{2}$ becomes progressively smaller. In fact, when $\theta\to\pi$ and $\sigma\to\infty$, $\mathbf{Z}_{1}+\mathbf{Z}_{2}$ becomes almost perpendicular to the individual noise vectors $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$. In this limit, $\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|_{\mathcal{H}}^{2}$ will approach $\gamma_{\epsilon,\delta}^{2}$. Note that in this case, the MSE of the sum is the same as if there were only one user; the noise from two users does not add up as it does in the LDP case. It is easy to generalize that even with $n$ users, the total MSE given by $\|\sum_{i=1}^{n}\mathbf{Z}_{i}\|^{2}_{\mathcal{H}}$ still approaches $\gamma_{\epsilon,\delta}^{2}$, and remains independent of $n$. In fact, this MSE is the same as in CDP! (See Appendix F-B for rigorous proofs).
Figure 4: Variation of the MSE with changing noise parameters $\sigma^{2}$ and $\rho$: Increasing $\|\mathbf{Z}_{i}\|_{\mathcal{H}}=\sigma\sqrt{d}$ and $\theta=\cos^{-1}\rho$ while maintaining the orthogonal distance between $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ at $\gamma_{\epsilon,\delta}$ for privacy, decreases $\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|_{\mathcal{H}}$ (the MSE).
@@
In essence, this example shows that for any given privacy parameters $\epsilon,\delta$, the optimized correlated Gaussian mechanism can achieve the same utility as CDP even without a trusted server. However, the correlated Gaussian scheme described above faces a significant shortcoming. If a user drops out, the residual noise of the remaining user/s cannot be canceled resulting in significantly high MSEs; this is because it uses noise with arbitrarily large variance (recall that the lowest possible MSE is achieved when $\|\mathbf{Z}_{i}\|_{\mathcal{H}}^{2}=E[||\mathbf{Z}_{i}||^{2}]\to\infty$). SecAgg—one of the main correlated privacy mechanisms used in practice—faces the same shortcoming. In SecAgg, users’ data is quantized to a finite field, and then a noise random variable that is uniformly distributed over the field elements is added to the quantized data and sent to the server. If a user drops out, this added noise cannot be canceled and the server input from the remaining users is statistically independent of the users’ data, which in effect results in a large MSE. SecAgg circumvents dropouts through additional rounds where remaining users’ share these non-canceled noise variables with the server enabling the server to cancel it. In fact, SecAgg (and its variants like GOPA [24]) is a complex protocol that requires more additional rounds to guard against further dropouts, data leaks and malicious behaviour in the previous rounds. Alternatively, this geometric interpretation offers a more straightforward method for addressing user dropouts, removing the necessity for additional communication rounds by appropriately adjusting the noise distribution. This gives rise to the CorDP-DME protocol introduced in this paper. In CorDP-DME we optimize the noise parameters—$\rho$ and $\sigma^{2}$—directly accounting for dropouts and colluding users (up to a threshold). As our analysis demonstrates, this approach significantly improves upon LDP in terms of MSE, while avoiding additional rounds of communication. In fact, in Section IV-A, we show that for any dropout threshold, CorDP-DME outperforms LDP with independent Gaussian noise.
IV Proposed Approach: CorDP-DME Protocol
In this section, we provide a comprehensive overview of CorDP-DME, along with a numerical example demonstrating the optimal noise parameters under different scenarios involving user dropouts and collusion. CorDP-DME operates in two phases: a data-independent offline phase, where correlated noise is generated (details in Section IV-D), and a data-dependent online phase. During the online phase, each user sends the encoded vectors from (1) to the server, which then computes the mean estimate using (2), completing the protocol. Similar to LDP-based DME, this approach remains simple and straightforward. Despite using correlated noise, CorDP-DME requires no additional communication rounds or computations to manage dropouts. The encoding and decoding processes in (1)-(2) are designed to handle dropouts and collusion seamlessly within the same round, while still ensuring privacy guarantees. Consequently, the primary focus of CorDP-DME is the design of the encoding and decoding functions, which is obtained by solving the optimization problem in (7).